# Solving Optimal Power Flow using a Variational Quantum Approach

### v1

@@ Abstract
The optimal power flow (OPF) is a large-scale optimization problem that is central in the operation of electric power systems. Although it can be posed as a nonconvex quadratically constrained quadratic program, the complexity of modern-day power grids raises scalability and optimality challenges. In this context, this work proposes a variational quantum paradigm for solving the OPF. We encode primal variables through the state of a parameterized quantum circuit (PQC), and dual variables through the probability mass function associated with a second PQC. The Lagrangian function can thus be expressed as scaled expectations of quantum observables. An OPF solution can be found by minimizing/maximizing the Lagrangian over the parameters of the first/second PQC. We pursue saddle points of the Lagrangian in a hybrid fashion. Gradients of the Lagrangian are estimated using the two PQCs, while PQC parameters are updated classically using a primal-dual method. We propose permuting primal variables so that OPF observables are expressed in a banded form, allowing them to be measured efficiently. Numerical tests on the IEEE 57-node power system using Pennylane’s simulator corroborate that the proposed doubly variational quantum framework can find high-quality OPF solutions. Although showcased for the OPF, this framework features a broader scope, including conic programs with numerous variables and constraints, problems defined over sparse graphs, and training quantum machine learning models to satisfy constraints.
1 Introduction
Optimal power flow (OPF) is a fundamental problem in the operation of electric power systems. It aims to determine the most economical, reliable, and secure schedule for generators to meet electric power demand while adhering to engineering limitations. Given the increasing load growth and amid a multi-faceted energy transition, grid operators must solve the OPF to optimally operate power systems of ever-increasing size and complexity at finer spatiotemporal scales.
OPF can be posed as an optimization problem over thousands of variables and constraints. If power flows need to be modeled precisely through the alternating-current (AC) power flow equations, the OPF can be formulated as a quadratically-constrained quadratic program (QCQP). Although this QCQP is not convex, the OPF can be relaxed to a convex semidefinite program (SDP) and solved to global optimality under certain conditions [1, 2]. Nonetheless, classical algorithms for solving SDPs [3], such as interior-point methods, may not scale well for large-scale OPF instances. Moreover, the SDP relaxation may not be always successful. In this context, and leveraging the success of variational quantum algorithms (VQAs) in other computational tasks, this work explores VQA alternatives to SDP solvers for addressing the OPF problem.
1.1 Contributions
The contribution of our paper is on three fronts:
c1) Modeling of the OPF using parameterized quantum circuits (PQC). The OPF can be posed as a non-convex QCQP. Both the number of variables and the number of constraints scale with the size of the power system. Given that modern power systems may consist of thousands of nodes, OPF is a large-scale optimization problem. To cope with the high dimensionality, we propose representing the OPF variables using two PQCs. Primal optimization variables are captured upon scaling the state of a primal PQC. Dual variables or Lagrange multipliers corresponding to constraints are captured by scaling the probability mass function (PMF) associated with a dual PQC. Due to this modeling, the Lagrangian function associated with the OPF can be expressed as an expectation over quantum observables operating on either or both PQCs. It is worth emphasizing that, although showcased for the OPF problem, the proposed modeling applies to other QCQP or more general problems that can be expressed in terms of expectations of quantum observables.
@@ 1.1 Contributions
c2) Optimizing a doubly parameterized Lagrangian function. To find approximate primal/dual OPF solutions, we propose a VQA seeking a saddle point of the associated Lagrangian function. In particular, we aim to minimize the Lagrangian function with respect to the parameters of the primal PQC and maximize it with respect to the parameters of the dual PQC. This is accomplished in a hybrid fashion. The two PQCs are used to measure gradients of the Lagrangian with respect to PQC parameters. A classical computer subsequently uses the gradient information to update PQC parameters iteratively based on the primal-dual method. The proposed primal-dual updates apply directly to a broader class of optimization problems that can be solved via VQAs. Moreover, the proposed method unlocks potential towards training quantum machine learning (QML) models to comply with constraints. Such functionality is essential for endowing QML models with safety and stability features.
@@
c3) Efficient measurement of OPF observables. Although OPF observables involve sparse Hermitian matrices, standard LCU-based decomposition methods do not yield qubit-efficient measurement protocols. To address this challenge, we utilize the structure of the OPF problem and the recently proposed method of extended Bell measurements (XBM) [4], to design a qubit-efficient protocol for measuring gradients of the Lagrangian function associated with the OPF. Specifically, the XBM method partitions the entries of an $N\times N$ Hermitian measurement matrix into groups, allowing entries within each group to be measured simultaneously using at most $\mathcal{O}(\log N)$ additional quantum gates. If a matrix can be decomposed into $\tilde{\mathcal{O}}(\log N)$ such groups, the XBM method is qubit-efficient. Two key points for successfully adopting the XBM method to the OPF are: i) The sparsity pattern (the positions of nonzero entries) of OPF observables is determined by the graph defined by the nodes and transmission lines of the power system, which is sparse; and ii) According to numerical tests, nodes in a power system can be permuted so that OPF observables can be partitioned in $\mathcal{O}((\log N)^{3})$ groups. Finding the node permutation that yields the number of groups is an NP-hard problem [5]. Instead, we adopt the reverse Cuthill-McKee (RCM) algorithm [6], a greedy heuristic algorithm of linear complexity $\mathcal{O}(N)$, to permute nodes so that OPF observables are banded. This is effective because banded matrices are known to feature a small number of groups [4]. Beyond the OPF setting, the devised workflow can be instrumental in effectively measuring quantum observables associated with sparse connectivity graphs arising in ML and optimization applications over natural or engineered networked systems.
1.2 Literature review
@@
Handling constrained optimization problems using VQAs. VQAs have emerged as a promising approach for solving optimization problems on current-day quantum computers. A well-studied exemplar of VQAs is the variational quantum eigensolver (VQE), which aims to find the smallest eigenvalue of a quantum observable by employing a PQC [13]. The quantum approximate optimization algorithm (QAOA) is a variant of VQE with a problem-dependent PQC and a diagonal observable targeting combinatorial problems [14]. While VQE and QAOA were originally devised for unconstrained problems, extensions to optimization problems with constraints have been proposed in [15, 16, 17, 18, 19, 20, 21, 22]. References [15, 19] convert constrained problems to unconstrained ones by penalizing constraint violations on the objective function. However, selecting a suitable penalty parameter is a nontrivial task. Unreasonably large values can cause ill-conditioning, while unreasonably small values can yield infeasible solutions. In [20], the QAOA’s ansatz is adapted to ensure that the output state stays within the feasible subspace; yet, this strategy applies only to binary problems with linear constraints. References [21] propose a hybrid method for solving SDPs. The idea is to approximate the matrix variable of the original SDP by a semidefinite matrix of smaller dimension. The resultant SDP of reduced dimension is solved classically using standard interior point-based methods. Quantum measurements are used to translate the cost and constraint functions of the original to the compressed SDP formulation. Nonetheless, the solution of the compressed SDP may be infeasible or suboptimal for the original SDP.
An alternative approach to cope with constrained problems is via Lagrangian duality. Reference [16] seeks a saddle point of the Lagrangian function with respect to the PQC parameters and the dual variables. To tackle problems with numerous constraints, reference [18] encodes dual variables through a parameterized quantum state, in addition to the primal variables. Both [16] and [18] solve the associated dual problem using the iterative method of dual decomposition. Nevertheless, each iteration of dual decomposition involves solving a VQE-type problem to optimality, which is computationally overwhelming. In [17], this drawback is circumvented by using the primal-dual update method, which only requires estimating gradients of the Lagrangian function. Nonetheless, reference [17] applies only to binary and linear programs. In this work, we combine [17] and [18] to deal with a doubly variational VQA formulation of the OPF.
Measuring quantum observables. Measuring quantum observables is key to the effective implementation of VQAs. A prominent measurement method decomposes a Hermitian matrix into a linear combination of unitaries (LCU), and measures the resultant unitary observables via the destructive swap test [23, 24]. This method is practically relevant only if the resultant unitaries can be implemented efficiently using quantum gates. This is the case, for example, if the unitaries can be expressed as Pauli strings (tensor products of Pauli matrices), and the number of Pauli strings scales polynomially with the number of qubits, $n$; see [25, 26, 27]. Unfortunately, a general matrix may involve $\mathcal{O}(4^{n})$ Pauli strings [28]. The number of Pauli strings grows exponentially with $n$, even for some structured sparse matrices, such as $k$-banded matrices that involve $\mathcal{O}(kn2^{n})$ Pauli strings [28]. One possible strategy to reduce the number of measurements is to group Pauli strings into sets of terms that pairwise commute. If two Hermitian matrices commute, they can be jointly diagonalized, and thus, measured simultaneously [29, pg. 43]. To this end, various grouping methods have been proposed, including qubit-wise commutativity grouping [30, 31], general commutativity grouping [32], and unitary partitioning [33, 34]. Nevertheless, partitioning Pauli strings to achieve the minimum number of groups is an NP-hard problem [32].
Another recent method for measuring observables involves extended Bell measurements [4]. This method decomposes the measurement matrix into matrices whose associated observables are simultaneously measurable using qubit-efficient rotation unitaries. Nevertheless, the number of induced observables may reach $2^{n}$ for a general sparse measurement matrix [4]. Our key idea is to permute nodes once and beforehand, so that OPF quantum observables can be measured efficiently.
1.3 Notation
@@
2 Optimal power flow as a QCQP
$\underaccent{\bar}{p}_{n}\leq p_{n}\leq\bar{p}_{n}\quad\text{and}\quad\underaccent{\bar}{q}_{n}\leq q_{n}\leq\bar{q}_{n}\quad\text{for all}\leavevmode\nobreak\ n\in\mathcal{N}.$
$\underaccent{\bar}{v}_{n}^{2}\leq\left|v_{n}\right|^{2}\leq\bar{v}_{n}^{2}\quad\text{for all}\leavevmode\nobreak\ n\in\mathcal{N}.$
Line current magnitudes are upper-limited by line capacities as
$\left|i_{\ell}\right|^{2}\leq\bar{i}_{\ell}^{2}\quad\text{for all}\leavevmode\nobreak\ \ell\in\mathcal{E}.$
Since all quantities mentioned above are quadratic in $\mathbf{v}$, the OPF can be posed as a quadratically constrained quadratic program (QCQP) over $\mathbf{v}$ as [2, 36]:
$P^{*}\coloneqq\min_{\mathbf{v}\in\mathbb{C}^{N}}$ $\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{0}\mathbf{v}$ (QCQP)
subject to (s.to) $\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{m}\mathbf{v}\leq b_{m},\quad\text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ m=1:M$
where $b_{m}$, for all $m$, is a given real-valued parameter determined by lower/upper limits and each $\mathbf{M}_{m}$ is a known Hermitian matrix. For example, if constraint $\underaccent{\bar}{p}_{n}\leq p_{n}$ is indexed as the $m$-th constraint, then $b_{m}=-\underaccent{\bar}{p}_{n}$. The OPF involves quadratic equality constraints, e.g., when $\underaccent{\bar}{p}_{n}=\bar{p}_{n}$ for a subset of nodes $n\in\mathcal{N}$. Each of these equality constraints corresponds to two inequality constraints in (QCQP). Overall, the number of OPF constraints can be several times the network size, such as $M\simeq 8N$.
The precise form of $\mathbf{M}_{0}$ and $\{(\mathbf{M}_{m},b_{m})\}_{m=1}^{M}$ is delineated in Appendix A.1. It is worth stressing that every matrix $\mathbf{M}_{m}$ is highly sparse. This is because every node of a power system is directly connected to only a few other nodes. The number of nonzero entries per matrix $\mathbf{M}_{m}$ is upper bounded by a small constant much smaller than $N$. The $(i,j)$-th off-diagonal entry across every matrix $\mathbf{M}_{m}$ is nonzero only if nodes $i$ and $j$ are connected through a transmission line, that is, if $(i,j)\in\mathcal{E}$. Consequently, the sparsity pattern (location of nonzero entries) for each $\mathbf{M}_{m}$ is determined by a small subset of $\mathcal{E}$. Therefore, the union of the sparsity patterns across all matrices $\mathbf{M}_{m}$ coincides with the sparsity pattern of the nodal admittance matrix $\mathbf{Y}$. This feature will be crucial for efficiently measuring the quantum observables involved in the OPF, as expounded in Section 6.
3 Primal-dual iterations for the OPF

### v2

# Solving Conic Programs over Sparse Graphs using a Variational Quantum Approach: The Case of the Optimal Power Flow

@@ Abstract
Conic programs arise broadly in physics, quantum information, machine learning, and engineering, many of which are defined over sparse graphs. Although such problems can be solved in polynomial time using classical interior-point solvers, the computational complexity scales unfavorably with graph size. In this context, this work proposes a variational quantum paradigm for solving conic programs, including quadratically constrained quadratic programs (QCQPs) and semidefinite programs (SDPs). We encode primal variables via the state of a parameterized quantum circuit (PQC), and dual variables via the probability mass function of a second PQC. The Lagrangian function can thus be expressed as scaled expectations of quantum observables. A primal-dual solution can be found by minimizing/maximizing the Lagrangian over the parameters of the first/second PQC. We pursue saddle points of the Lagrangian in a hybrid fashion. Gradients of the Lagrangian are estimated using the two PQCs, while PQC parameters are updated classically using a primal-dual method. We propose permuting the primal variables so that related observables are expressed in a banded form, enabling efficient measurement. The proposed framework is applied to the OPF problem, a large-scale optimization problem central to the operation of electric power systems. Numerical tests on the IEEE 57-node power system using Pennylane’s simulator corroborate that the proposed doubly variational quantum framework can find high-quality OPF solutions. Although showcased for the OPF, this framework features a broader scope, including conic programs with numerous variables and constraints, problems defined over sparse graphs, and training quantum machine learning models to satisfy constraints.
preprint: APS/123-QED
I Introduction
Conic programs defined over sparse graphs are ubiquitous in physics, quantum information, machine learning, and engineering. Many of these problems naturally take the form of semidefinite programs (SDPs) or nonconvex quadratically-constrained quadratic programs (QCQPs) that can be lifted to SDPs. In many of these formulations, the number of constraints scales with the graph size, and thus, although each constraint involves a small subset of variables, the Lagrangian function contains numerous terms. Representative examples include finding the ground-state energy of an Ising spin system [4], computing graph-theoretic quantities such as the Lovász theta number [42], verifying deep neural network properties [53], and networked optimization in engineering, such as the optimal power flow problem [3, 38]. Although classical algorithms for solving SDPs, such as interior-point methods [8], are guaranteed to find a globally optimal solution, their computational cost may scale unfavorably with the problem dimension.
Variational quantum algorithms (VQAs) have emerged as a versatile and promising paradigm for leveraging noisy intermediate-scale quantum (NISQ) processors to solve complex optimization problems. By employing parameterized quantum circuits (PQCs) optimized via classical routines, these hybrid quantum-classical algorithms are designed to comply with the constraints of current quantum hardware. Nonetheless, their application to constrained optimization remains underdeveloped. Existing methodologies are mostly limited to problems with binary variables and/or linear constraints, leaving a gap for the more general yet powerful class of conic programs, such as QCQPs and SDPs. Recently, Reference [51] used two PQCs to encode both primal and dual variables (Lagrange multipliers) for solving SDPs with many constraints. Nevertheless, it assumes efficient measurement of observables and relies on a dual-decomposition method that entails a nested variational optimization.
I.1 Contributions
This work proposes a novel VQA framework for solving large-scale conic programs with numerous constraints defined over sparsely connected graphs. The contributions are on three fronts:
@@ I.1 Contributions
c1) Optimizing a doubly parameterized Lagrangian function. Building on the modeling proposed in [51], where primal variables are captured by the state of a primal PQC and the dual variables are captured through the probability mass function (PMF) associated with a dual PQC, we develop a VQA seeking a saddle point of the associated Lagrangian function. We aim to minimize the Lagrangian function with respect to the parameters of the primal PQC and maximize it with respect to the parameters of the dual PQC. This is accomplished in a hybrid fashion. First, the two PQCs are used to measure gradients of the Lagrangian with respect to the PQC parameters. A classical computer subsequently uses gradient information to iteratively update PQC parameters using a primal-dual method. In contrast to the dual decomposition [51, 24], which requires solving an inner variational optimization problem to optimality at each iteration, our method performs a single gradient descent step on the primal variables and a single gradient ascent step on the dual variables per iteration. The proposed primal-dual updates apply directly to a broader class of optimization problems that can be solved via VQAs. Moreover, the proposed method enables the training of quantum machine learning (QML) models that satisfy constraints. Such functionality is essential for endowing QML models with safety and stability features.
@@
c2) Efficient measurement of graph-induced observables. Although the concerned observables are defined over sparse Hermitian matrices, standard LCU (linear combination of unitaries)-based decomposition methods do not yield qubit-efficient measurement protocols. To address this challenge, we utilize the graph structure of the problem instance and the extended Bell measurements (XBM) [33], to design a qubit-efficient protocol for measuring gradients of the Lagrangian function associated with the QCQP/SDP. Specifically, the XBM method partitions the entries of an $N\times N$ Hermitian measurement matrix into groups, allowing entries within each group to be measured simultaneously using at most $\mathcal{O}(\log N)$ additional quantum gates. If a matrix can be decomposed into $\tilde{\mathcal{O}}(\log N)$ such groups, the XBM method is qubit-efficient. Hence, we propose to permute the graph’s nodes so that the related observables are banded. Finding the node permutation that yields the optimal number of groups is an NP-hard problem [18]. Instead, we adopt the reverse Cuthill-McKee (RCM) algorithm [14], a greedy heuristic algorithm of linear complexity $\mathcal{O}(N)$. This is effective because banded matrices are known to feature a small number of XBM groups [33]. The devised workflow can be instrumental in effectively measuring quantum observables associated with sparse connectivity graphs arising in applications over natural or engineered networked systems, including the representative examples discussed earlier.
c3) Application to the Optimal Power Flow (OPF) problem. We illustrate the efficacy of the proposed framework on the OPF problem. OPF can be posed as an optimization problem over thousands of variables and constraints. If power flows need to be modeled precisely through the alternating-current (AC) power flow equations, the OPF can be formulated as a QCQP. Although this QCQP is not convex, the OPF can be relaxed to a convex SDP and solved to global optimality under certain conditions [3, 38]. Nonetheless, the complexity of modern-day electric power grids raises scalability and optimality challenges. In this context, VQA offers a promising alternative, provided that the required quantum measurements can be performed efficiently. Two key points for successfully adopting the XBM method to the OPF are: i) The sparsity pattern (the positions of nonzero entries) of OPF observables is determined by the graph defined by the nodes and transmission lines of the power system, which is sparse; and ii) According to numerical tests, nodes in a power system can be permuted via the RCM algorithm so that OPF observables can be partitioned in $\mathcal{O}((\log N)^{3})$ groups.
I.2 Literature review
Handling constrained optimization problems using VQAs. A well-studied exemplar of VQAs is the variational quantum eigensolver (VQE), which aims to find the smallest eigenvalue of a quantum observable by employing a PQC [52]. The quantum approximate optimization algorithm (QAOA) is a variant of VQE with a problem-dependent PQC and a diagonal observable targeting combinatorial problems [17]. While VQE and QAOA were originally devised for unconstrained problems, extensions to optimization problems with constraints have been proposed in [43, 24, 39, 51, 11, 25, 6, 7]. References [43, 11] convert constrained problems to unconstrained ones by penalizing constraint violations on the objective function. However, selecting a suitable penalty parameter is a nontrivial task. Unreasonably large values can cause ill-conditioning, while unreasonably small values can yield infeasible solutions. In [25], the QAOA’s ansatz is adapted to ensure that the output state stays within the feasible subspace; yet, this strategy applies only to binary problems with linear constraints. References [6] propose a hybrid method for solving SDPs. The idea is to approximate the matrix variable of the original SDP by a semidefinite matrix of smaller dimension. The resultant SDP of reduced dimension is solved classically using standard interior point-based methods. Quantum measurements are used to translate the cost and constraint functions of the original to the compressed SDP formulation. Nonetheless, the solution of the compressed SDP may be infeasible or suboptimal for the original SDP.
An alternative approach to cope with constrained problems is via Lagrangian duality. Reference [24] seeks a saddle point of the Lagrangian function with respect to the PQC parameters and the dual variables. To tackle problems with numerous constraints, reference [51] encodes dual variables through a parameterized quantum state, in addition to the primal variables. Both [24] and [51] solve the associated dual problem using the iterative method of dual decomposition. Nevertheless, each iteration of dual decomposition involves solving a VQE-type problem to optimality, which is computationally overwhelming. In [39], this drawback is circumvented by using the primal-dual update method, which only requires estimating gradients of the Lagrangian function. Nonetheless, reference [39] applies only to binary and linear programs. In this work, we extend [39] and [51] to deal with a doubly variational VQA formulation of QCQP/SDPs, which we showcase via the OPF problem.
Measuring quantum observables. Measuring quantum observables is key to the effective implementation of VQAs. A prominent measurement method decomposes a Hermitian matrix into a linear combination of unitaries (LCU), and measures the resultant unitary observables via the destructive swap test [59, 21]. This method is practically relevant only if the resultant unitaries can be implemented efficiently using quantum gates. This is the case, for example, if the unitaries can be expressed as Pauli strings (tensor products of Pauli matrices), and the number of Pauli strings scales polynomially with the number of qubits, $n$; see [46, 12, 49]. Unfortunately, a general matrix may involve $\mathcal{O}(4^{n})$ Pauli strings [36]. The number of Pauli strings grows exponentially with $n$, even for some structured sparse matrices, such as $k$-banded matrices that involve $\mathcal{O}(kn2^{n})$ Pauli strings [36]. One possible strategy to reduce the number of measurements is to group Pauli strings into sets of terms that pairwise commute. If two Hermitian matrices commute, they can be jointly diagonalized, and thus, measured simultaneously [60, pg. 43]. To this end, various grouping methods have been proposed, including qubit-wise commutativity grouping [62, 63], general commutativity grouping [23], and unitary partitioning [29, 64]. Nevertheless, partitioning Pauli strings to achieve the minimum number of groups is an NP-hard problem [23].
Another recent method for measuring observables involves extended Bell measurements [33]. This method decomposes the measurement matrix into matrices whose associated observables are simultaneously measurable using qubit-efficient rotation unitaries. Nevertheless, the number of induced observables may reach $2^{n}$ for a general sparse measurement matrix [33]. Our key idea is to permute nodes of the underlying graph once and beforehand, so that the associated quantum observables can be measured efficiently.
@@ I.2 Literature review
I.3 Notation
@@
II Primal-dual iterations for QCQP
We focus on optimization problems defined over sparse graphs, often arising from physical models as nonconvex QCQPs or SDPs. The latter may appear either as native formulations or as convex relaxations of the QCQPs. Prototypical examples of such problems include:
  • Ising ground-state energy: Minimizing a quadratic Hamiltonian $H(\mathbf{s})=\sum_{i<j}J_{ij}s_{i}s_{j}+\sum_{i}h_{i}s_{i}$ over spin configurations $\mathbf{s}\in\{\pm 1\}^{N}$ can be expressed as a nonconvex QCQP. The coupling matrix $\mathbf{J}$ is typically sparse, reflecting local interactions in physical lattices or graphical models. The resulting QCQP admits an SDP relaxation obtained by lifting the quadratic $\pm 1$ constraints $s_{i}^{2}=1$; see [4].
  • Lovász theta number: For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $N=|\mathcal{V}|$ vertices, the Lovász theta number can be computed as the optimal value of an SDP over an $N\times N$ matrix variable:
    $\vartheta(\mathcal{G}):=\max_{\mathbf{X}\succeq 0}\penalty 10000\$$\penalty 10000\ \mathbf{1}^{\top}\mathbf{X}\mathbf{1}$
    subject to (s.to) $\penalty 10000\ \operatorname{Tr}(\mathbf{X})=1,\quad$
    $\penalty 10000\ \mathbf{X}_{i,j}=0\ \forall(i,j)\in\mathcal{E},$
    where $\mathbf{1}$ is the all-ones vector. Each edge $(i,j)\in\mathcal{E}$ yields one constraint, so the constraint matrices inherit the sparsity pattern of $\mathcal{G}$. While commonly presented as an SDP, the Lovász theta number admits an equivalent QCQP formulation [42].
  • Neural network verification: Certifying the robustness of neural networks (NN) against adversarial perturbations entails bounding the worst‑case margin between classes over a permissible input set. In networks using the rectified linear unit (ReLU) as activation function, the input-output relationship of the ReLU $z=\max(0,x)$ admits an exact reformulation including a quadratic constraint, namely $z\geq 0$, $z\geq x$, and z(zx)=0z(z-x)=0 [53]. Based on this reformulation, the NN verification task can be cast as a nonconvex QCQP over the high-dimensional vector of neuron activations. The constraint matrices are highly sparse and structured, reflecting the layered connectivity of the underlying computational graph [53].
  • Optimal power flow: The AC OPF problem minimizes the electricity generation cost to meet load demand while satisfying grid engineering constraints. The problem can be posed as a nonconvex QCQP over nodal voltages, where the quadratic constraints represent Kirchhoff’s laws and operational limits. The sparsity pattern is governed by the power grid topology through transmission lines, as each grid node is directly connected to only a few other nodes [3].
To streamline the presentation, we focus on the QCQP formulation and discuss its extension to the SDP setting in Section III. We consider the QCQP:
$P^{*}\coloneqq\min_{\mathbf{x}\in\mathbb{C}^{N}}\penalty 10000\$$\penalty 10000\ \mathbf{x}^{\dagger}\mathbf{M}_{0}\mathbf{x}$ (QCQP)
s.to $\penalty 10000\ \mathbf{x}^{\dagger}\mathbf{M}_{m}\mathbf{x}\leq b_{m},\quad\text{for}\penalty 10000\ \penalty 10000\ m=1:M$
where $\{\mathbf{M}_{m}\in\mathbb{C}^{N\times N}\}_{m=0}^{M}$ are sparse Hermitian matrices determined by the underlying problem graph, and the number of constraints $M$ typically scales with the network size $N$.